Imagine you started your own taxi business in Toronto. The average customer takes approximately 3 rides and spends $70 a month. For each ride, 50% of the fare goes to the driver, and $2 goes to the administrative and maintenance costs. Assuming 36% of riders find alternative travel arrangements and don’t use the taxi company the following year and the discount rate is 3% per year. What is the CLV?


The answer is $865.54

Can you show me how to get that answer please

1
Expert's answer
2021-11-02T03:50:02-0400

CLV margin is the difference between the revenue you receive from a customer and all of the costs associated with that customer in a given timeframe. To get CLV, it is then multiplied by your retention rate and divided by one plus your average discount rate minus your retention rate.

"CLV=Revenue*Time *(1+retention\/rate)"

"Revenue=70"

Timespan; since 36% find alternative means then in a year "=0.64*12"

Time Span "=7.68"

Rate"=[1+(0.03-0.64)" "=0.39"

Then the retention should be "=1-0.39 = 0.61"

Therefore the CLV should be;

"CLV=70*7.68*1.61"

"CLV=865.54"